EWhen a block of metal of mass 0.11kg and specific heat capacity 400J/kg/k is Heated to 100°c and quickly transferred to a calorimeter containing 0.20kg of a liquid at 10°c the resulting temperature of 18°c. On repeating the experiment with 0.40kg of liquid in the same container and at the same initial temperature of 10°c the resulting temperature is 14.5°c. Calculate a. The specific heat capacity of the liquid b. The thermal capacity of the container.


1
Expert's answer
2021-11-14T17:10:14-0500

The energy acquired by the block is released to the water and container in both scenarios. For the two experiments:


"Q_b+Q_l+Q_c=0,\\
c_bm_b(t_b-t_1)+c_lm_{l1}(t_l-t_1)+C_c(t_l-t_1)=0,\\
c_bm_b(t_b-t_2)+c_lm_{l2}(t_l-t_2)+C_c(t_l-t_2)=0.\\\space\\
c_l=1925\text{ j/(kg}·\text{K)},\\
C_c=66\text{ J/K}."